OpenStudy (firechild17):

Which of the following is the correct graph of the solution to the inequality −8 greater than or equal to −5x + 2 > −38? number line with an open dot at 2 with shading to the left and another open dot at 8 with shading to the right. number line with an open dot at 8 and at 2 with shading in between. number line with a closed dot on 2 and an open dot on 8 and shading in between. number line with closed dots at 2 and 8 with shading between the 2 and the 8.

OpenStudy (firechild17):

@Frozen15

OpenStudy (jackthegreatest):

\[-8\ge-5x+2>-38\]

OpenStudy (jackthegreatest):

what value of x will make -5x + 2=-38

OpenStudy (jackthegreatest):

@firechild17

OpenStudy (firechild17):

so its B Im here I was reading lol

OpenStudy (jackthegreatest):

yep b

OpenStudy (firechild17):

thanks

OpenStudy (jackthegreatest):

wait nope

OpenStudy (jackthegreatest):

the dot should not be open for 2

OpenStudy (firechild17):

huh?

OpenStudy (firechild17):

so its C?

OpenStudy (jackthegreatest):

because \[-8\ge-5x + 2\]

OpenStudy (firechild17):

ok

OpenStudy (jackthegreatest):

the greater than or EQUAL to includes 2

OpenStudy (jackthegreatest):

yes C

OpenStudy (firechild17):

ok